# Impact of the valence band on Rydberg excitons in cuprous oxide quantum wells

### v1

@@ IV Conclusion and outlook
In this article, we have considered the complete valence band structure but have still used some approximations and simplifications, which shall be addressed in future work. In the Hamiltonian (8) we have used the same Luttinger parameters as in the bulk material [10]. This may be a good approximation for sufficiently broad QWs ($L\gtrsim 5$ nm), however, the Luttinger parameters probably need to be adjusted for narrow QWs based on density functional theory (DFT) calculations [53] of the electronic band structure of the thin films [54]. In our calculations we have also assumed the same dielectric constant $\varepsilon$ for both the crystal and the substrate. The dielectric contrast of a more realistic setup can be taken into account via the Rytova-Keldysh potential [42, 43] similar to what has been done for the hydrogenlike model of excitons in QWs [28]. Furthermore, we have restricted our computations to a configuration in which the QW is aligned perpendicular to the $[001]$ axis of the crystal, and excitons are invariant under symmetry transformations of the $D_{4\mathrm{h}}$ symmetry group. A more general orientation between the QW and the crystal can be considered, however, this will further reduce the symmetries of the problem. For the computation of spectra and oscillator strengths we have focused on bound states of the yellow exciton series. In future work, this can be generalized to yellow and green resonance states above threshold by application of the complex-coordinate-rotation method as in Ref. [36]. This will also enable the search for BICs [37]. The consideration of all these effects will certainly allow for detailed comparisons with future experimental spectra.

### v2

@@ IV Conclusion and outlook
In this article, we have considered the complete valence band structure but have still used some approximations and simplifications, which shall be addressed in future work. In the Hamiltonian (10) we have used the same Luttinger parameters as in the bulk material [10]. This may be a good approximation for sufficiently broad QWs ($L\gtrsim 5$ nm), however, the Luttinger parameters probably need to be adjusted for narrow QWs based on density functional theory (DFT) calculations [62] of the electronic band structure of the thin films [63]. In our calculations we have also assumed the same dielectric constant $\varepsilon$ for both the crystal and the substrate. The dielectric contrast of a more realistic setup can be taken into account via the Rytova-Keldysh potential [46, 47] similar to what has been done for the hydrogenlike model of excitons in QWs [31]. Furthermore, we have restricted our computations to a configuration in which the QW is aligned perpendicular to the $[001]$ axis of the crystal, and excitons are invariant under symmetry transformations of the $D_{4\mathrm{h}}$ symmetry group. A more general orientation between the QW and the crystal can be considered, however, this will further reduce the symmetries of the problem. For the computation of spectra and oscillator strengths we have focused on bound states of the yellow exciton series. In future work, this can be generalized to yellow and green resonance states above threshold by application of the complex-coordinate-rotation method as in Ref. [39]. This will also enable the search for BICs [40]. The obtained wave functions will make it possible to estimate the interband dipole transition matrix elements to investigate the usage of the confined Rydberg excitons for quantum sensing. The consideration of all these effects and radiative properties will certainly allow for detailed comparisons with future experimental spectra.
Acknowledgements.
Stefan Scheel and Pavel A. Belov acknowledge financial support by the National Research Council Canada, grant No. QSP-129-1.